Could you please solve this question in a short time ?

QUESTION Obtain phase margin of the system giver as disturbance load PI controller Reference (s+0.5) 14 output S st4
Answer #1

given, disturbance output sto.5 5 5+4 5to.5 sty itis+u) M >output Sto 5 9+8 S Sto.5 :) ( ) open loop transfer function GCs).finding again 1G(s)HG) = 1 at Wgc w2+0.52 = w2+82 45 w²60.52 ww²+g2 squaring both sides 16(weroisy 16w2 +4= wyt 64 W² wy 48w²Pwgc -0.79 X 180 Í deg --45:38 Owgc -H5.30 180 t phage magin Owge 180 45-3 0 134.69 phaje margin